OpenStudy (anonymous):

Find the range of the function \(f(x) = x^\frac{1}{x}\)

OpenStudy (sidsiddhartha):

i think [-infinity,infinity)

OpenStudy (anonymous):

Domain: (-∞,1) U (1,∞) Range: (-∞,0) U (0,∞)

OpenStudy (shubhamsrg):

y= x^(1/x) log y = x/logx y'/y = (1/logx) - 1 y' = y( (1/logx) -1) since y>0, 1/logx - 1 =0 (for finding the extremity) this yield x=e hence maxima or minima occurs at x=e or at (e,e^1/e)

OpenStudy (shubhamsrg):

It can be easily seen that it'd be a point of maxima since 1) it is the only extremity of the graph 2) 2<e and 4>e ; (2^1/2) = (4^1/4) Hence the graph has the same heights at both x=2 and 4. It must have come up till e^1/e and then again fallen down.

OpenStudy (anonymous):

Hmmm............ I am not sure if you have the second line of your first comment right. I think it should be y= x^(1/x) log y = (1/x) log x log y = (log x) / x

OpenStudy (anonymous):

y= x^(1/x) logy = (logx)/x y' / y = (1-logx) / x^2 y' = y [ (1-logx) / x^2 ] Set y' = 0, it is known that y > 0, y' =0 y [ (1-logx) / x^2 ] =0 (1-logx) / x^2 =0 x = e y = e^(1/e) when x=e. Then applies your argument here :D Thanks for the idea! I forgot we can use calculus here!!

OpenStudy (shubhamsrg):

oops :P i miscalculated the derivative :P

OpenStudy (anonymous):

I guess the first problem was you made a mistake when you took log on both sides?!

OpenStudy (shubhamsrg):

Right. Sorry about that. That was silly of me

OpenStudy (anonymous):

It's alright! Human errs :) Thanks a lot again for your kind help! :D

OpenStudy (shubhamsrg):

Glad if I could help ^^